OpenStudy (anonymous):

x+y=3 substitute y=x-1

OpenStudy (anonymous):

just replace y with x - 1, solve for x and then y after that

OpenStudy (anonymous):

x+x-1=3

OpenStudy (anonymous):

mhmm go on

OpenStudy (anonymous):

i dnt know what to do next

OpenStudy (anonymous):

add 1 to both sides, then simply

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

what is x + x?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

thats x * x

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

2x+4

OpenStudy (anonymous):

x+x-1=3, add one, u get x + x = 4

OpenStudy (anonymous):

and since x + x is 2 x, wut is x when 2x = 4?